OpenStudy (anonymous):

would someone please help me find the limit of (e^-x + 2cos(3x) as x approaches inifinity? I'm stuck on the math. I came up with 0 as an answer, but I think I'm wrong. Would someone help me break it down step by step to find my error? Thank you.

OpenStudy (anonymous):

\[\lim_{x\to \infty} e^{-x}=0\] but \(\lim_{x\to \infty}\cos(x)\) does not exist, as cosine is periodic

OpenStudy (anonymous):

cosine oscillates between -1 and 1 infinitely often

OpenStudy (anonymous):

ok..I see what you are saying...so in this case because the two limits are not equal, then this limit does not exist...is that what you are saying?